Using the discriminant or otherwise, find the value of k for which the equation 3kx^2 + 4x + 2=0 has equal roots
Answer by htmentor(1343)

The quadratic ax^2 + bx + c will have only one root if b^2 - 4ac = 0
4^2 = 4*3k*2
16 = 24k
k = 2/3
Check:
3(2/3)x^2 + 4x + 2 = 0
2x^2 + 4x + 2 = 0
x^2 + 2x + 1 = 0
(x+1)^2 = 0
x = -1